If the work required to speed a car from 10 km/h to 20 km/h is 5.0 x 103 J, what would be the work required to increase its speed from 20 km/h to 30 km/h.


1
Expert's answer
2022-05-18T15:28:11-0400
"W_{12}=\frac{m}{2}(v_2^2-v_1^2)"

"W_{34}=\frac{m}{2}(v_4^2-v_3^2)"

"W_{34}=W_{12}\frac{v_4^2-v_3^2}{v_2^2-v_1^2}"

"W_{34}=5.0*10^3*\frac{30^2-20^2}{20^2-10^2}=8.3*10^3\:\rm J"